# Whose Emotions and Moral Sentiments Do Language Models Reflect?

### v1

@@ 3.2 Representativeness of Affect in Steered Prompting
Figure 3 and 6 (in Appendix) provides insights into how steering instruction-tuned LMs and base LMs respectively to adopt a liberal ($g_{l}$) or conservative ($g_{c}$) persona impacts affective alignment measured by emotions. Figure 7 (in Appendix) shows the affective alignment measured by moral foundations. The directionality of triangle symbols shows the nature of steering: left for liberal steering and right for conservative steering. The circles show the models’ baselines, i.e. the default alignment which are identical to the circles in Figure 2 and Figure 5 (in Appendix).
Is steering effective? We expect that a model’s affective alignment with an ideological group after liberal steering and conservative steering should differ; otherwise, we deem that the steering is ineffective. In Figure 3, it is evident that steering is effective for most instruction-tuned LMs, as indicated by the left-facing and right-facing triangles of the same color positioned apart from each other. However, such failure cases happen for almost all base LMs, as indicated by the the left-facing and right-facing triangles of the same color positioned extremely close to each other or even overlapping in Figure 6 (in Appendix). This observation demonstrates that instruction-tuning and RLHF make LMs more steerable. We do not exclude the possibility that the failure cases for base LMs are caused by the specific prompts we used to steer the base LMs (Table 3 in Appendix), but we leave how to craft better prompts to steer base LMs for future work. In the regard, in the following analysis related to steering, we only focus on instruction-tuned models.
@@
Do the models exhibit strong affective alignment after steering? Although steering enhances affective alignment for most instruction-tuned LMs (Figure  3), the alignment of LMs to either ideological group is still lower than the partisan alignment baseline. Notably, the more sophisticated model gpt-3.5, even after steering, is least aligned with both partisan perspectives.
@@ 4 Related Work
LMs trained on extensive datasets of human language from the Internet, are capable of simulating realistic discourse. To ensure that LMs generate text consistent with human values and ethical principles, many recent works have investigated the human-LM alignment. Popular frameworks include reinforcement learning with human feedback (RLHF) or AI feedback (RLAIF) Ouyang et al. (2022); Glaese et al. (2022); Bai et al. (2022). To measure alignment Santurkar et al. (2023) compared LMs’ opinions with human responses in public opinion polls among various demographic groups and found substantial misalignment. Durmus et al. (2023) expanded the study of alignment to a global scale using cross-national surveys and discovered LMs’ inclination towards certain countries like USA, as well as unwanted cultural stereotypes. Zhao et al. (2023) proposed steering language models to better fit individual groups. Simmons (2022) measured LMs’ moral biases associated with political groups in the United States when responding to different moral scenarios; however, they only evaluate the models’ moral responses based on a general statistical finding from previous works that “liberals rely primarily on individualizing foundations while conservatives make more balanced appeals to all 5 foundations”. In contrast, our work evaluate the models against affect distributions observed from real-world human-generated texts on a topic basis.
@@
Feng et al. (2023) discovered that pretrained LMs do exhibit political biases, propagating them into downstream tasks. In terms of adapting LMs to simulate human opinions, Argyle et al. (2023) showed that GPT-3 can mimic respondents in extensive, nationally-representative opinion surveys. Other researchers have finetuned LMs to learn the political views of different partisan communities to study polarization Jiang et al. (2022); He et al. (2024). To evaluate news feed algorithms, Törnberg et al. (2023) created multiple LM personas from election data to simulate conversations on social media platforms.
@@ 5 Conclusion
Affective Classifier Accuracies The classifiers used for emotions and moralities are not perfect. However, our method depends on comparing the emotion and morality distributions between the real-world and model-generated tweets. This comparative approach mitigates the impact of potential classifier inaccuracies, as the same classifier is applied consistently across both corpora. Since we are primarily looking at differences, rather than absolute values of emotions in the data, we believe we are justified in using the imperfect classifiers to reveal differences in affective alignment. Nevertheless, we have endeavored to utilize the most advanced models currently available for accurately measuring emotions and moral foundations in the sociopolitical domain. The performance of both models has been validated on a variety of social media data Rao et al. (2023); Guo et al. (2023a); Chochlakis et al. (2023), and proposing methods to achieve the new state-of-the-art on emotion and morality detection is out of the scope of this work.
@@
Ethics Statement and Broader Impact
@@
Ethical Impact and Data Use. Our work utilizes publicly available data from social media, specifically Twitter, which poses potential privacy concerns. We have ensured that all Twitter data used in our study has been accessed in compliance with Twitter’s data use policies and that individual privacy has been respected, with no attempt to de-anonymize or reveal personally identifiable information. The dataset consists of tweets related to COVID-19 and Roe v. Wade, which are topics of public interest and social importance. In handling this data, we were careful to maintain the anonymity of the users and to treat the content with the utmost respect, given the sensitive nature of the topics.
Potential Applications and Broader Impacts. The potential applications of our work range from enhancing the empathetic capabilities of LLMs to ensuring that AI systems can understand and respect diverse perspectives. While these are positive outcomes, we recognize the possibility of misuse, such as the reinforcement of biases or the manipulation of public discourse. To mitigate such risks, we recommend that any application of our findings be accompanied by rigorous fairness and bias assessment protocols.
Appendix A Appendix
A.1 Online Sociopolitical Discourse Data
@@
We compile two datasets on sociopolitical discourse on Twitter: COVID-19 Tweets and Roe v. Wade Tweets. They cover a wide range of fine-grained topics, including emotionally divisive topics. To assess the affect alignment, we identify important issues discussed in the Twitter datasets using a semi-supervised method described in Rao et al. (2023). This method harvests and selects from Wikipedia the relevant and distinctive keywords for each issue, and detect the issues in each tweet using the presence of these keywords and phrases. An issue, such as “masking” in COVID-19 tweets, can still be broad and too general. In order to obtain a fine-grained span of topics, we use GPT-4 to cluster the keywords in each issue into sub-topics, such as “mask mandates and policies” and “mask health concerns”. We manually validated the clustering results. Each tweet can be associated with multiple issues and sub-topics.
COVID-19 Tweets
@@
Roe v. Wade Tweets
A.2 Experimental Setup
On each topic, we obtain 2,000 generations from a model.
@@ A.2 Experimental Setup
Figure 5: Default affect alignment $S$ of different LMs with both ideological groups – liberals ($g_{l}$) and conservatives ($g_{c}$), measured by moral foundations. For each LM, the alignment is averaged over that on different topics detected within the dataset, with the means shown by circles and the standard deviations shown by errors bars. Base LMs and instruction-tuned LMs are separated by the black horizontal dashed line. The alignment between the two ideological groups (above the red horizontal dashed line) themselves are measured as a baseline.
@@
Figure 6: Steered affective alignment $S$ of different LMs with both ideological groups – liberals ($g_{l}$) and conservatives ($g_{c}$), measured by emotions, for base LMs. Left-facing triangles represent the models by liberal steered prompting; right-facing triangles represent the models by conservative steered prompting; circles with no filling colors represent the models by default. For each LM, the alignment is averaged over that on different topics detected within the dataset. The alignment between the two ideological groups (above the red horizontal dashed line) themselves are measured as a baseline.
@@
Figure 7: Steered affect alignment $S$ of different LMs with ideological groups – liberals ($g_{l}$) and conservatives ($g_{c}$), measured by moral foundations. Left-facing triangles represent the models by liberal steered prompting; right-facing triangles represent the models by conservative steered prompting; circles with no filling colors represent the models by default. For each LM, the alignment is averaged over that on different topics detected within the dataset. Base LMs and instruction-tuned LMs are separated by the black horizontal dashed line. The alignment between the two ideological groups (above the red horizontal dashed line) themselves are measured as a baseline.

### v2

@@ 3.2 Representativeness of Affect in Steered Prompting
Figure 3 provides insights into how steering instruction-tuned LMs to adopt a liberal ($g_{l}$) or conservative ($g_{c}$) persona impacts affective alignment measured by emotions. Please refer to Appendix D.3 for that of base LMs measured by emotions, and that of all LMs measured by moral sentiments. The directionality of triangle symbols shows the nature of steering: left for liberal steering and right for conservative steering. The circles show the models’ baselines, i.e. the default alignment which are identical to the circles in Figure 2.
Is steering effective? We expect that a model’s affective alignment with an ideological group after liberal steering and conservative steering should differ; otherwise, we deem that the steering is ineffective. In Figure 3, it is evident that steering is effective for most instruction-tuned LMs, as indicated by the left-facing and right-facing triangles of the same color positioned apart from each other. However, such failure cases happen for almost all base LMs, as indicated by the the left-facing and right-facing triangles of the same color positioned extremely close to each other or even overlapping (Figure 6 in Appendix D.3). This observation demonstrates that instruction-tuning and RLHF make LMs more steerable. We do not exclude the possibility that the failure cases for base LMs are caused by the specific prompts we used to steer the base LMs, but we leave how to craft better prompts to steer base LMs for future work. In the regard, in the following analysis related to steering, we only focus on instruction-tuned models.
@@
Do the models exhibit strong affective alignment after steering? Although steering enhances affective alignment for most instruction-tuned LMs, the alignment of nearly all LMs to either ideological group is still lower than the partisan alignment baseline. Notably, the more sophisticated model gpt-3.5, even after steering, is least aligned with both partisan perspectives.
@@ 4 Related Work
LMs trained on extensive datasets of human language from the Internet, are capable of simulating realistic discourse. To ensure that LMs generate text consistent with human values and ethical principles, many recent works have investigated the human-LM alignment. Popular frameworks include reinforcement learning with human feedback (RLHF) or AI feedback (RLAIF) Ouyang et al. (2022); Glaese et al. (2022); Bai et al. (2022). To measure alignment Santurkar et al. (2023) compared LMs’ opinions with human responses in public opinion polls among various demographic groups and found substantial misalignment. Durmus et al. (2023) expanded the study of alignment to a global scale using cross-national surveys and discovered LMs’ inclination towards certain countries like USA, as well as unwanted cultural stereotypes. Simmons (2022) measured LMs’ moral biases associated with political groups in the United States when responding to different moral scenarios; however, they only evaluate the models’ moral responses based on a general statistical finding from previous works that “liberals rely primarily on individualizing foundations while conservatives make more balanced appeals to all 5 foundations”. Abdulhai et al. (2023) measured moral foundations of LLMs using the Moral Foundations Questionnaire (MFQ) questionnaire and compared the results of LLMs to humans. In contrast, our work evaluate the models against affect distributions observed from real-world human-generated texts on a topic basis. The affect people/models express in open-ended texts is likely to be different from how they answer the close-ended questionnaire, and our framework can capture the fluidity and complexity of human affective responses in a way that structured questionnaires might not fully encompass.
@@
Feng et al. (2023) discovered that pretrained LMs do exhibit political biases, propagating them into downstream tasks. In terms of adapting LMs to simulate human opinions, Argyle et al. (2023) showed that GPT-3 can mimic respondents in extensive, nationally-representative opinion surveys. Other researchers have finetuned LMs to learn the political views of different partisan communities to study polarization Jiang et al. (2022); He et al. (2024). To evaluate news feed algorithms, Törnberg et al. (2023) created multiple LM personas from election data to simulate conversations on social media platforms. Chen et al. (2024) found that most LLMs show a left-leaning stance on a wide range of issues and LLMs can be easily manipulated by instruction tuning.
@@ 5 Conclusion
Acknowledgements
This project has been funded, in part, by DARPA under contract HR00112290106 and HR00112290106. We appreciate the constructive advice and suggestions from the anonymous reviewers.
@@ Acknowledgements
Affective Classifier Accuracies The classifiers used for emotions and moralities are not perfect. However, our method depends on comparing the emotion and morality distributions between the real-world and model-generated tweets. This comparative approach mitigates the impact of potential classifier inaccuracies, as the same classifier is applied consistently across both corpora. Since we are primarily looking at differences, rather than absolute values of emotions in the data, we believe we are justified in using the imperfect classifiers to reveal differences in affective alignment. Nevertheless, we have endeavored to utilize the most advanced models currently available for accurately measuring emotions and moral foundations in the sociopolitical domain. The performance of both models has been validated on a variety of social media data Rao et al. (2023b); Guo et al. (2023a); Chochlakis et al. (2023), and proposing methods to achieve the new state-of-the-art on emotion and morality detection is out of the scope of this work.
@@
Ethics Statement
@@
Our work utilizes publicly available data from social media, specifically Twitter, which poses potential privacy concerns. We have ensured that all Twitter data used in our study has been accessed in compliance with Twitter’s data use policies and that individual privacy has been respected, with no attempt to de-anonymize or reveal personally identifiable information. The dataset consists of tweets related to COVID-19 and Roe v. Wade, which are topics of public interest and social importance. In handling this data, we were careful to maintain the anonymity of the users and to treat the content with the utmost respect, given the sensitive nature of the topics.
Appendix A Broader Impact
Implications of affective alignment.
Introducing affective alignment, our paper bridges a gap left by prior research focused predominantly on positional alignment. The impact of affective alignment of LMs is crucial in the following contexts.
In mental health applications, an LM’s ability to align affectively with users can provide support and improve therapeutic outcomes. In educational settings, affective alignment in LMs can foster a conducive learning environment, adapting to students’ emotional states to enhance engagement and comprehension. In political discourse, affective alignment is key to fostering constructive debates and reducing polarization; for instance, during election campaigns, LMs that can align affectively with the emotional and moral sentiments of different voter bases can facilitate more resonant and effective communication.
However, there are also scenarios where affective alignment is undesirable. For example, In high-stakes negotiation or diplomacy settings, excessive affective alignment might hinder the ability to maintain a firm stance or negotiate effectively. In legal or judicial contexts, an overly empathetic LM could bias the presentation of facts or arguments, potentially affecting impartiality. Similarly, in news reporting, high affective alignment might lead to biased news coverage, undermining journalistic neutrality.
Therefore, it is important to achieve a balanced affective alignment that enhances positive interactions and outcomes without compromising on objectivity and fairness. Recognizing these subtleties can guide the ethical and effective deployment of LMs, ensuring they serve as beneficial tools in society rather than exacerbating existing divides or biases.
Implications of LM representativeness of affect across different demographics.
The representativeness of LMs in capturing affect across different demographics is crucial for ensuring that AI systems communicate in ways that are emotionally and culturally resonant. However, the representativeness of LMs in reflecting affect across demographics should be contextually calibrated, not merely equalized. For example, an LM used in a global social media platform should accurately reflect the emotional and moral nuances of its worldwide user base, avoiding over-representation of any single group’s affective norms. In contrast, an LM in localized service applications, like community-based mental health support, should be tuned to the specific emotional and cultural characteristics of that community. Balancing these representational needs requires a strategic approach to developing LMs that are both inclusive and contextually aware.
Implications of the framework to measure affective alignment.
Our proposed framework offers a comprehensive tool for assessing how well LMs resonate with human emotions and morals across various demographics and subjects. This versatility facilitates broader research applications, enabling researchers and developers to evaluate and enhance LM designs for cultural sensitivity, inclusiveness, and ethical alignment with human values, thus paving the way for more responsible AI innovations.
Implications of our findings on the affective alignment of LMs with both political groups.
Our analysis reveals that LMs display a notable liberal bias, especially in contexts like COVID-19 discussions, and exhibit an overall affective misalignment with political groups, surpassing the existing partisan divide in the U.S. This indicates a systemic inclination within LMs. These observations underscore the challenge of LMs achieving strong affective alignment with humans and equal affective representations.
In conclusion, our research not only contributes to the academic understanding of LMs but also serves as a pivotal guide for developing AI that is emotionally intelligent, morally considerate, and socially representative.
Appendix B Online Sociopolitical Discourse Data
We compile two datasets on sociopolitical discourse on Twitter: COVID-19 Tweets and Roe v. Wade Tweets. They cover a wide range of fine-grained topics, including emotionally divisive topics. Our selection of COVID-19 and Roe vs. Wade Twitter datasets was based on the following factors:
  • COVID-19 has global significance, affecting diverse aspects of life and eliciting a wide range of emotional and moral responses, making it ideal for studying affective alignment. Roe v. Wade, representing a longstanding and polarizing issue in U.S. politics, provides a rich dataset to explore deeply entrenched moral and emotional viewpoints, allowing for a detailed analysis of language models’ alignment with complex ideological positions.
  • These datasets are publicly accessible, which is crucial for ensuring the transparency and reproducibility of our research.
  • The method for estimating user partisanship Rao et al. (2021) is particularly effective for these datasets. This is because the tweets in these datasets frequently include URLs to news articles, which serve as reliable indicators of the users’ political leanings. We’ll add these considerations to the paper.
@@ Appendix B Online Sociopolitical Discourse Data
To assess the affect alignment, we identify important issues discussed in the Twitter datasets using a semi-supervised method described in Rao et al. (2023b). This method harvests and selects from Wikipedia the relevant and distinctive keywords for each issue, and detect the issues in each tweet using the presence of these keywords and phrases. An issue, such as “masking” in COVID-19 tweets, can still be broad and too general. In order to obtain a fine-grained span of topics, we use GPT-4 to cluster the keywords in each issue into sub-topics, such as “mask mandates and policies” and “mask health concerns”. We manually validated the clustering results. Each tweet can be associated with multiple issues and sub-topics.
B.1 COVID-19 Tweets
@@
Appendix C Proximity between Emotions
The polar coordinates of the 11 emotions (anger, anticipation, disgust, fear, joy, love, optimism, pessimism, sadness, surprise, and trust) are $-\frac{1}{2}\pi$, $-\frac{1}{4}\pi$, $-\frac{3}{4}\pi$, $\frac{1}{2}\pi$, 00, $\frac{1}{8}\pi$, $-\frac{1}{8}\pi$, $\frac{7}{8}\pi$, $\pi$, $\frac{3}{4}\pi$, and $\frac{1}{4}\pi$ respectively. For emotions that are not the Plutchik-8 emotions, we aggregate the coordinates of their neighboring emotions. The emotion proximity matrix is shown in Table 3.
anger anticipation disgust fear joy love optimism pessimism sadness surprise trust
anger 1 0.75 0.75 0 0.5 0.375 0.625 0.375 0.5 0.25 0.25
anticipation 1 0.5 0.25 0.75 0.625 0.875 0.125 0.25 0 0.5
disgust 1 0.25 0.25 0.125 0.375 0.625 0.75 0.5 0
fear 1 0.5 0.625 0.375 0.625 0.5 0.75 0.75
joy 1 0.875 0.875 0.125 0 0.25 0.75
love 1 0.75 0.25 0.125 0.375 0.875
optimism 1 0 0.125 0.125 0.625
pessimism 1 0.875 0.875 0.375
sadness 1 0.75 0.25
surprise 1 0.5
trust 1
Table 3: Proximity scores between different emotions. We only show the upper triangle as it is a symmetric matrix.
Appendix D Experiments
D.1 Experimental Setup
On each topic, we obtain 2,000 generations from a model, using the prompt templates shown in Table 4.
@@ D.1 Experimental Setup
D.2 Representativeness of Affect under Default Prompting
Figure 5 shows the affective alignment of various LMs with liberals ($g_{l}$) and conservatives ($g_{c}$) by default prompting in the two datasets measured by moral sentiments.
@@ D.2 Representativeness of Affect under Default Prompting
Figure 5: Default affect alignment $S$ of different LMs with both ideological groups – liberals ($g_{l}$) and conservatives ($g_{c}$), measured by moral foundations. * indicates that the alignment of the liberal steered model with both ideological groups are significantly different at $p<0.05$. For each LM, the alignment is averaged over that on different topics detected within the dataset, with the means shown by circles and the standard deviations shown by errors bars. Base LMs and instruction-tuned LMs are separated by the black horizontal dashed line. The alignment between the two ideological groups (above the red horizontal dashed line) themselves are measured as a baseline.
D.3 Representativeness of Affect under Steered Prompting
Figure 6 provides insights into how steering instruction-tuned LMs to adopt a liberal ($g_{l}$) or conservative ($g_{c}$) persona impacts affective alignment measured by moral sentiments. Figure 7 shows the affective alignment under steered prompting measured by moral foundations.
@@ D.3 Representativeness of Affect under Steered Prompting
Figure 6: Steered affective alignment $S$ of different base LMs with both ideological groups – liberals ($g_{l}$) and conservatives ($g_{c}$), measured by emotions . Left-facing triangles represent the models by liberal steered prompting; right-facing triangles represent the models by conservative steered prompting; circles with no filling colors represent the models by default. * indicates that the alignment of the liberal steered model with both ideological groups are significantly different at $p<0.05$; ^ indicates that for the conservative steered model. For each LM, the alignment is averaged over that on different topics detected within the dataset. The alignment between the two ideological groups (above the red horizontal dashed line) themselves are measured as a baseline.
@@
Figure 7: Steered affect alignment $S$ of different LMs with ideological groups – liberals ($g_{l}$) and conservatives ($g_{c}$), measured by moral foundations. Left-facing triangles represent the models by liberal steered prompting; right-facing triangles represent the models by conservative steered prompting; circles with no filling colors represent the models by default. * indicates that the alignment of the liberal steered model with both ideological groups are significantly different at $p<0.05$; ^ indicates that for the conservative steered model. For each LM, the alignment is averaged over that on different topics detected within the dataset. Base LMs and instruction-tuned LMs are separated by the black horizontal dashed line. The alignment between the two ideological groups (above the red horizontal dashed line) themselves are measured as a baseline.
D.4 Topic-Level Analysis
Figure 8 shows emotion and moral foundation distributions of tweets from two LMs – gpt-3.5 and llama-2-7b-chat – and humans from both ideological groups, on the topic “fetal rights debate in abortion” from the Roe v. Wade Tweets.